OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -8) and a directrix at y = 8.

hero (hero):

If you have two points, the focus \((x_1, y_1)\) and the directrix \((x_2, y_2)\), then you can insert both points in to the formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\)

OpenStudy (anonymous):

Here are the answer choices: y =-1/32x^2 y^2 = -32x y^2 = -8x y =-/8x^2

OpenStudy (anonymous):

So how I go about solving the problem now?

hero (hero):

In this case, the focus is \((0,-8)\) The directrix is \((0,8)\) If you inserted the given points in to the equation, you'd have: \((x - 0)^2 + (y - (-8))^2 = (x - x)^2 + (y - 8)^2\)

hero (hero):

Then you'd simplify that to just \(x^2 + (y + 8)^2 = (y - 8)^2\) The next step is to expand the binomials and solve for y

OpenStudy (anonymous):

Let me try that:3

OpenStudy (anonymous):

Is the answer y =-1/32x^2

hero (hero):

Show the work you did to arrive at that answer.